8. A financial analyst has found out that policyholders are four times as likely to file two
claims as to file four claims. If the number of claims filed has a posson distribution,
what is the variance of the number of claims filed.

The lifetime of a machine is continuous on the interval (0, 40) with probability density
function f, where f(t) is proportional to (t + 10)) 2
, and t is the lifetime in years.
Calculate the probability that the lifetime of the machine part is less than 10 years.
Hint: Show that f(t) is legitimate and find the proportionality constant
1
Expert's answer
2020-05-11T13:07:45-0400

1. We want to find "\lambda" for the Poisson RV from "P(X=2)=4P(X=4)"


"P(X=k)={e^{-\lambda}\lambda^k\over k!}"

"P(X=2)={e^{-\lambda}\lambda^2\over 2!}=4P(X=4)=4\cdot{e^{-\lambda}\lambda^4\over 4!}"

"\lambda^2={4!\over4\cdot2!}=3"

"V(X)=\sigma^2=\lambda=\sqrt{3}"


2.


"f(t) = \begin{cases}
 c(t+10)^2 & t\in[0,40] \\
 0 &otherwise
\end{cases}"

"\displaystyle\int_{-\infin}^{\infin}f(t)dt=1"

"\displaystyle\int_{0}^{40}c(t+10)^2dt=c\big[{(t+10)^3\over 3}\big]\begin{matrix}
 40 \\
 0
\end{matrix}="

"={c\over 3}((40+10)^3-(0+10)^3)={124000\over 3}c=1"

"c={3\over 124000}"

"P(t<10)=\displaystyle\int_{0}^{10}{3\over 124000}(t+10)^2dt="

"={3\over 124000}\big[{(t+10)^3\over 3}\big]\begin{matrix}
 10 \\
 0
\end{matrix}={1\over 124000}((10+10)^3-(0+10)^3)="

"={7\over 124}\approx0.056452"